Light of wavelength 550 nm (in a vacuum) passes through a diamond with a refractive index of 2.42.
What is the wavelength of the light while it is in the diamond?
1
Expert's answer
2020-05-29T09:51:39-0400

Wavelength of light moving in vacuum is "\lambda=550nm=550\times10^{-9}m" .

Refractive Index of Diamond is "\mu=2.42" .

Let, frequency of light is "\nu" and wavelength of light in Diamond is "\lambda_1".


Since, refractive index is the ratio of speed of light in vacuum to the speed of light in any other medium,we get

"\mu=\frac{c}{v}"

But we also know that frequency of light is independent of medium i.e frequency of light doesn't change, thus

"\mu=\frac{\nu \lambda}{\nu \lambda_1}\\
\implies \mu=\frac{\lambda}{ \lambda_1}\\"

Hence, from the above formula we get,


"\lambda_1=\frac{\lambda}{\mu}\\
\implies \lambda_1=\frac{550\times 10^{-9}}{2.42}=227.27nm"